OpenStudy (anonymous):

halp plz Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem.

OpenStudy (anonymous):

halp cri

OpenStudy (texaschic101):

they purchase flowers for 50 cents they sell flowers for 50 cents + 40% of purchase price... 50 + 0.40(50) = 50 + 20 = 70 cents so this means they make a profit of (70 - 50).. 20 cents for every flower they sell. .20f = 500 -- " f " stands for flowers f = 500/0.20 f = 2500 check... bought 2500 flowers for 50 cents each = 0.50 * 2500 = $1250 sold 2500 flowers for 70 cents each = 0.70 * 2500 = $ 1750 for a profit of (1750 - 1250) = $ 500